Magnetic flux may be varied by
1
Expert's answer
2020-03-31T09:19:53-0400

Magnetic flux


"\Phi=B\cdot A\cdot \cos\alpha"


You can change "\Phi" by changing "B" and "A" or "\alpha" .

Also, "B" can be altered by changing the current.


If "B=const" and "A=const" then "\Phi" changes from "B\cdot A (\alpha=0°)" to "-B\cdot A (\alpha=180°)".